# DexterCap: An Affordable and Automated System for Capturing Dexterous Hand-Object Manipulation

### v1

@@ 3.4.1. Hand Model Parameterization
To ensure biomechanical plausibility, we define local coordinate systems for the joints of the MANO model following (Spurr et al., 2020), as shown in Figure 4. We parameterize hand pose using anatomically informed degrees of freedom (DoF): the thumb’s metacarpophalangeal (MCP) and proximal interphalangeal (PIP) joints are assigned 3 DoF each, while its distal interphalangeal (DIP) joint is modeled with 1 DoF. For the other fingers, each MCP joint is also assigned 3 DoF, and both PIP and DIP joints are assigned 1 DoF. This parameterization yields a reduced pose vector $\bm{\phi}\in\mathbb{R}^{27}$, which is mapped in a differentiable manner to the full MANO pose space $\theta\in\mathbb{R}^{45}$. The global hand translation is represented as $\bm{t}\in\mathbb{R}^{3}$, and the global orientation $\bm{o}$ is modeled using a 6D rotation representation (Zhou et al., 2019).
Before capturing data for a new subject, we first estimate the hand shape parameters $\beta\in\mathbb{R}^{10}$ by fitting the MANO model to a 3D scan of the subject’s hand. This scan does not need to be high-precision and can be acquired using an affordable 3d scanner. In our experiments, we employ a low-cost Structure Sensor 1 depth camera mounted on an iPhone and operated with its official scanning application under default settings, which yields a coarse mesh (around 6k vertices) without any additional post-processing. The fitting is performed by minimizing the Chamfer Distance between the model and the scan, with optional supervision from direct finger length measurements. We optimize $\beta$ using gradient descent. Once estimated, $\beta$ is fixed for that subject and used in all subsequent motion capture sessions.
@@
Figure 4. Local joint coordinate systems defined for the MANO model.
Figure 5. Manually defined submeshes of the MANO hand, with each color corresponding to a specific finger segment.
3.4.2. Calibration
Solving hand motion begins with a calibration process to associate the predefined markers with the MANO model, similar to prior non-rigid registration work(White et al., 2007; Halimi et al., 2022). Formally, this association can be formulated as
(1) $\hat{\bm{x}}_{i}=M_{i}(\bm{t},\bm{o},\bm{\phi},{b})=b_{i}^{1}v_{i}^{1}+b_{i}^{2}v_{i}^{2}+b_{i}^{3}v_{i}^{3}$
@@ 3.4.2. Calibration
Figure 6. Hand-object interaction in the DexterHand dataset.
@@ 3.4.3. Solving Motion
(3) $E_{\text{reg}}=\sum_{k\in N_{\text{DoF}}}\|\text{clip}(\bm{\phi}_{k},\bm{\phi}_{k,\text{low}},\bm{\phi}_{k,\text{high}})-\bm{\phi}_{k}\|_{2}^{2}$
where the prior limits $\bm{\phi}_{\text{low}}$ and $\bm{\phi}_{\text{high}}$ are manually defined to cover the natural range of joint angles.
The total objective function is then
(4) $E=E_{\text{point}}+\lambda_{\text{reg}}E_{\text{reg}}$
We optimize this objective function using Adam (Kingma and Ba, 2014) optimizer with a learning rate of 0.002 for 400 epochs.
@@
3.5. Object Pose Estimation
@@
Figure 7. Object pose estimation.
Table 1. Overview of selected motion capture sequences in the DexterHand dataset. Shape dimensions and penetrations are in centimeters. Notations: $x$, $y$, $z$: length, width, and height of cuboid; $h$: height; $d$: diameter; $D$: outer diameter (ring); $b$: side length of a triangular prism.
@@
For sequences with object interaction, we estimate the object’s 6-DoF pose (position and orientation) using known rigid 3D mesh models. We employ the Kabsch algorithm (Kabsch, 1976) for all frames to align observed object markers $\bm{X}_{\text{obs}}$ with canonical marker positions $\bm{X}_{\text{canonical}}$. This algorithm determines the optimal rotation $\bm{R}_{\text{obj}}$ and translation $\bm{t}_{\text{obj}}$ by minimizing the following objective function:
(5) $L_{\text{obj}}=\|\bm{X}_{\text{obs}}-(\bm{R}_{\text{obj}}\bm{X}_{\text{canonical}}+\bm{t}_{\text{obj}})\|_{2}^{2}$
@@ 4. DexterHand Dataset
5. Evaluation
5.1. System Setup
@@
5.2. Image Processing and Marker Recognition
@@
Corner detection. Our CornerNet (heatmap-based) achieves 94.7% precision, 81.6% recall, and 87.7% F1 score at the image level, where a corner is considered correctly detected if it lies within 5 pixels of the ground truth. In contrast, our implementation of the coordinate-regression model in (Chen et al., 2021a) reaches 98.3% accuracy but only 46.3% precision, yielding many false positives on patches without markers.
@@
Notably, our edge detection model enables the edge-first assembly strategy to propose candidate quadrilaterals for block identification. Compared to the exhaustive quadrilateral validation strategy of (Chen et al., 2021a), the edge-first assembly reduces the average number of candidates from 5550 quads per frame to 83 blocks (via  707 edges), greatly improving efficiency and overall quality.
To show this, we reimplemented the method from (Chen et al., 2021a). Although the model achieves block-level recognition accuracy exceeding 99%, the extremely large number of quadrilateral candidates, most of which are negatives, substantially lowers its performance, resulting in only 55.0% precision and 23.6% recall.
@@
5.3. Hand-Object Reconstruction
@@
5.4. Motion Quality and Plausibility
@@
Benchmark against other motion capture methods. We benchmark our dataset and pipeline against several public datasets: GRAB (Taheri et al., 2020), ARCTIC (Fan et al., 2023), and HUMOTO (Lu et al., 2025a), which are captured using commercial marker-based systems or data gloves. For these datasets, we compute all metrics from the released trajectories using the same implementation as for our data. We also consider the vision-only methods HaMeR (Pavlakos et al., 2024) as an additional baseline. In particular, HaMeR is a single-view, vision-based method. We apply it to our captures by running it independently on our camera view. Since its performance can degrade on our grayscale videos, and because the markers attached to the fingers may alter the hand appearance relative to its training distribution, we further include the GigaHands (Fu et al., 2025) dataset as an additional baseline. GigaHands is a multi-view extension of HaMeR, and we compute the same metrics on its released trajectories.
@@
Table 2. Benchmark comparison with vision-only and commercial systems. Bold values indicate best performance in each metric.
@@
6. Future Work and Conclusions
@@ Appendix A DANCE CARD FOR RUBIK’S CUBE
Table 3. Dance Card for $\times2$$\times$2 Rubik’s Cube Manipulation
@@ 15 Rotate the down face clockwise and counter-clockwise by 2 loops (8 times) $(D4)2(D^{\prime}4)2$
16 Do an algorithm [^1] This algorithm is also known as the ”Sexy Move”.4 times $(RUR^{\prime}U^{\prime})6$
@@ B.1. System Overview
B.2. Marker Configuration and Reference Model
@@
B.3. Coplanarity-Based Rotation Detection
@@
(7) $\text{Score}(\mathcal{P})=\sigma_{\min}(\bm{P}-\bar{\bm{p}})$
@@
B.4. 122-Block Decomposition and Registration
@@
B.5. Relative Rotation Estimation
@@
(9) $\theta=\text{sign}(\bm{n}^{T}\cdot\bm{\omega})\|\bm{\omega}\|$
where $\bm{\omega}$ is the rotation vector from $\bm{R}_{\text{rel}}=\bm{R}_{2}\bm{R}_{1}^{-1}$, and $\bm{n}$ is the rotation axis vector transformed by $\bm{R}_{1}$.
B.6. Implementation Details
@@ Appendix C Model Architectures
Table 4. Architecture of CornerNet.
@@
As summarized in Table 4, CornerNet adopts a U-Net-based (Ronneberger et al., 2015) encoder-decoder architecture with four resolution levels and symmetric skip connections.
EdgeNet is built upon a ResNet-34 (He et al., 2016) backbone, adapted to single-channel inputs, and followed by a lightweight fully connected prediction head.
@@ Appendix E Motion Quality Metrics
MSNR. A higher Motion Signal-to-Noise Ratio (MSNR) indicates smoother motion. Following HUMOTO (Lu et al., 2025a), we compute
(10) $\mathrm{MSNR}=10\log_{10}\left(\frac{\mathbb{E}[\hat{v}^{2}]}{\mathbb{E}[|v-\hat{v}|^{2}]}\right).$

### v2

@@ 3.4.1 Hand Model Parameterization
To ensure biomechanical plausibility, we define local coordinate systems for the joints of the MANO model following [spurr2020weakly], as shown in Figure 3. We parameterize hand pose using anatomically informed degrees of freedom (DoF): the thumb’s metacarpophalangeal (MCP) and proximal interphalangeal (PIP) joints are assigned 3 DoF each, while its distal interphalangeal (DIP) joint is modeled with 1 DoF. For the other fingers, each MCP joint is also assigned 3 DoF, and both PIP and DIP joints are assigned 1 DoF. This parameterization yields a reduced pose vector $\bm{\phi}\in\mathbb{R}^{27}$, which is mapped in a differentiable manner to the full MANO pose space $\theta\in\mathbb{R}^{45}$. The global hand translation is represented as $\bm{t}\in\mathbb{R}^{3}$, and the global orientation $\bm{o}$ is modeled using a 6D rotation representation [zhou2019continuity].
Before capturing data for a new subject, we first estimate the hand shape parameters $\beta\in\mathbb{R}^{10}$ by fitting the MANO model to a 3D scan of the subject’s hand. This scan does not need to be high-precision and can be acquired using an affordable 3d scanner. In our experiments, we employed a low-cost Structure Sensor 1 depth camera mounted on an iPhone and operated with its official scanning application under default settings, which yields a coarse mesh (around 6k vertices) without any additional post-processing. The fitting is performed by minimizing the Chamfer Distance between the model and the scan, with optional supervision from direct finger length measurements. We optimize $\beta$ using gradient descent. Once estimated, $\beta$ is fixed for that subject and used in all subsequent motion capture sessions.
@@
Figure 3: Local joint coordinate systems defined for the MANO model.
Figure 4: Manually defined submeshes of the MANO hand, with each color corresponding to a specific finger segment.
3.4.2 Calibration
Solving hand motion begins with a calibration process to associate the predefined markers with the MANO model, similar to prior non-rigid registration work[white07capturing, halimi22pattern]. Formally, this association can be formulated as
$\hat{\bm{x}}_{i}=M_{i}(\bm{t},\bm{o},\bm{\phi},{b})=b_{i}^{1}v_{i}^{1}+b_{i}^{2}v_{i}^{2}+b_{i}^{3}v_{i}^{3}$ (1)
@@ 3.4.2 Calibration
Figure 5: Hand-object interaction in the DexterHand dataset.
@@ 3.4.3 Solving Motion
$E_{\text{reg}}=\sum_{k\in N_{\text{DoF}}}\|\text{clip}(\bm{\phi}_{k},\bm{\phi}_{k,\text{low}},\bm{\phi}_{k,\text{high}})-\bm{\phi}_{k}\|_{2}^{2}$ (3)
where the prior limits $\bm{\phi}_{\text{low}}$ and $\bm{\phi}_{\text{high}}$ are manually defined to cover the natural range of joint angles. The total objective function is then
$E=E_{\text{point}}+\lambda_{\text{reg}}E_{\text{reg}}$ (4)
We optimize this objective function using Adam [Kingma2014AdamAM] optimizer with a learning rate of 0.002 for 400 epochs.
@@
3.5 Object Pose Estimation
@@
Figure 6: Object pose estimation.
Table 1: Overview of selected motion capture sequences in the DexterHand dataset. Shape dimensions and penetrations are in centimeters. Notations: $x$, $y$, $z$: length, width, and height of cuboid; $h$: height; $d$: diameter; $D$: outer diameter (ring); $b$: side length of a triangular prism.
@@
For sequences with object interaction, we estimate the object’s 6-DoF pose (position and orientation) using known rigid 3D mesh models. We employ the Kabsch algorithm [kabsch1976solution] for all frames to align observed object markers $\bm{X}_{\text{obs}}$ with canonical marker positions $\bm{X}_{\text{canonical}}$. This algorithm determines the optimal rotation $\bm{R}_{\text{obj}}$ and translation $\bm{t}_{\text{obj}}$ by minimizing the following objective function:
$L_{\text{obj}}=\|\bm{X}_{\text{obs}}-(\bm{R}_{\text{obj}}\bm{X}_{\text{canonical}}+\bm{t}_{\text{obj}})\|_{2}^{2}$ (5)
@@ 4 DexterHand Dataset
5 Evaluation
5.1 System Setup
@@
5.2 Image Processing and Marker Recognition
@@
Corner detection. Our CornerNet (heatmap-based) achieves 94.7% precision, 81.6% recall, and 87.7% F1 score at the image level, where a corner is considered correctly detected if it lies within 5 pixels of the ground truth. In contrast, our implementation of the coordinate-regression model in [chen2021capturing] reaches 98.3% accuracy but only 46.3% precision, yielding many false positives on patches without markers.
@@
Notably, our edge detection model enables the edge-first assembly strategy to propose candidate quadrilaterals for block identification. Compared to the exhaustive quadrilateral validation strategy of [chen2021capturing], the edge-first assembly reduces the average number of candidates from 5550 quads per frame to 83 blocks (via  707 edges), greatly improving efficiency and overall quality.
To show this, we reimplemented the method from [chen2021capturing]. Although the model achieves block-level recognition accuracy exceeding 99%, the extremely large number of quadrilateral candidates, most of which are negatives, substantially lowers its performance, resulting in only 55.0% precision and 23.6% recall.
@@
5.3 Hand-Object Reconstruction
@@
5.4 Motion Quality and Plausibility
@@
Benchmark against other motion capture methods. We benchmark our dataset and pipeline against several public datasets: GRAB [taheri2020grab], ARCTIC [fan2023arctic], and HUMOTO [lu2025humoto], which are captured using commercial marker-based systems or data gloves. For these datasets, we compute all metrics from the released trajectories using the same implementation as for our data. We also consider the vision-only methods HaMeR [pavlakos2024reconstructing] as an additional baseline. In particular, HaMeR is a single-view, vision-based method. We apply it to our captures by running it independently on one of our camera views. Since its performance can degrade on our grayscale videos, and because the markers attached to the fingers may alter the hand appearance relative to its training distribution, we further include the GigaHands [fu2025gigahands] dataset as an additional baseline. GigaHands is a multi-view extension of HaMeR, and we compute the same metrics on its released trajectories.
@@
Table 2: Benchmark comparison with vision-only and commercial systems. Bold values indicate best performance in each metric.
@@
6 Future Work and Conclusions
@@ Appendix A DANCE CARD FOR RUBIK’S CUBE
Table 3: Dance Card for $\times2$$\times$2 Rubik’s Cube Manipulation
@@ 15 Rotate the down face clockwise and counter-clockwise by 2 loops (8 times) $(D4)2(D^{\prime}4)2$
16 Do an algorithm [^1] This algorithm is also known as the "Sexy Move".4 times $(RUR^{\prime}U^{\prime})6$
@@ B.1 System Overview
B.2 Marker Configuration and Reference Model
@@
B.3 Coplanarity-Based Rotation Detection
@@
$\text{Score}(\mathcal{P})=\sigma_{\min}(\bm{P}-\bar{\bm{p}})$ (7)
@@
B.4 122-Block Decomposition and Registration
@@
B.5 Relative Rotation Estimation
@@
$\theta=\text{sign}(\bm{n}^{T}\cdot\boldsymbol{\omega})\|\boldsymbol{\omega}\|$ (9)
where $\boldsymbol{\omega}$ is the rotation vector from $\bm{R}_{\text{rel}}=\bm{R}_{2}\bm{R}_{1}^{-1}$, and $\bm{n}$ is the rotation axis vector transformed by $\bm{R}_{1}$.
B.6 Implementation Details
@@ Appendix C Model Architectures
Table 4: Architecture of CornerNet.
@@
As summarized in Table 4, CornerNet adopts a U-Net-based [ronneberger2015unet] encoder-decoder architecture with four resolution levels and symmetric skip connections.
EdgeNet is built upon a ResNet-34 [he2015resnet] backbone, adapted to single-channel inputs, and followed by a lightweight fully connected prediction head.
@@ Appendix E Motion Quality Metrics
MSNR. A higher Motion Signal-to-Noise Ratio (MSNR) indicates smoother motion. Following HUMOTO [lu2025humoto], we compute
$\mathrm{MSNR}=10\log_{10}\left(\frac{\mathbb{E}[\hat{v}^{2}]}{\mathbb{E}[|v-\hat{v}|^{2}]}\right).$ (10)